# Quantum Signatures of Chaos from Free Probability

### v1

@@ 5 Asymptotic Freeness in the High-Spin Mixed-Field Ising Model
$[X^{(s)}_{j},Y^{(s)}_{k}]=i\frac{\sqrt{3}}{\sqrt{s(s+1)}}\delta_{jk}Z^{(s)}_{j},$ (59)
@@ 5.1 Approximate emergence of asymptotic freeness
In the spin-1 case, the fundamental operators of the theory, such as $X^{(s)}_{i}$, and their time-evolved counterparts, $X^{(s)}_{i}(t)$, have spectra characterized by a generalized Bernoulli distribution with eigenvalues ±$\pm\sqrt{3/2},0$. For convenience, we rescale these operators by a factor of $\sqrt{2/3}$, ensuring their spectra become ±$\pm 1,0$. As shown in Section 3.2 (Example 2), the free probability prediction for the spectrum of the sum of two free variables with generalized Bernoulli distribution with eigenvalues ±$\pm 1,0$ is given by Eq. (39). Figure 21 shows the spectrum of eigenvalues of the operator $\sqrt{2/3}(Z_{1}(0)+Z_{1}(t))$ with increasing time for 100 realizations of the Hamiltonian (57) at the chaotic point. The distribution predicted by free probability theory (39) becomes evident around $t=5$.
[…]
@@ 5.2 Fluctuations
In this subsection, we investigate the fluctuations in the spectrum of the sum of spin-1 operators of the form $A+B(t)$. In the presence of chaotic dynamics, the operators $A$ and $B(t)$ are expected to become asymptotically free, with the distribution of their sum given by Eq. (39). The results are presented in Figures 22, 23, and 24. In all cases, we consider 100 realizations of the Hamiltonian (57) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$, and $t=50$.
@@
Figure 22: Spectral statistics of the operator $\sqrt{2/3}(Z_{1}^{(s)}+Z_{6}^{(s)}(t))$ based on 100 realizations of the Hamiltonian (57) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$ and $t=50$. The solid lines represent (a) the free probability prediction and (b) the Wigner $r$-parameter distribution for $\nu=2$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=1.999\pm 0.004$. The average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% on each edge.
@@
Figure 23: Spectral statistics of the operator $\sqrt{2/3}(Y_{1}^{(s)}+Y_{6}^{(s)}(t))$ based on 100 realizations of the Hamiltonian (57) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$ and $t=50$. The solid lines represent (a) the free probability prediction and (b) the Wigner $r$-parameter distribution for $\nu=2$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=2.000\pm 0.004$. The average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% on each edge.
@@
Figure 24: Spectral statistics of the operator $\sqrt{2/3}(X_{1}^{(s)}+X_{6}^{(s)}(t))$ based on 100 realizations of the Hamiltonian (57) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$ and $t=50$. The solid lines represent (a) the free probability prediction and (b) the Wigner $r$-parameter distribution for $\nu=2$. The $r$-parameter distribution was fitted using the form (83), yielding the Wigner parameter $\nu=2.000\pm 0.004$. The average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% on each edge.
@@ 6 Asymptotic Freeness in the Sachdev–Ye–Kitaev Model
$\psi_{i}=\frac{1}{\sqrt{2}}\begin{cases}\bigotimes_{k=1}^{\nu}\sigma_{1}&\text{if }i=1,\\[10.0pt] \left(\bigotimes_{k=1}^{\nu-\left\lfloor\frac{i}{2}\right\rfloor}\sigma_{1}\right)\otimes\sigma_{2}\otimes\left(\bigotimes_{k=1}^{\left\lfloor\frac{i}{2}\right\rfloor-1}\sigma_{0}\right)&\text{if }2\leq i\leq 2\nu\text{ and }i\text{ is even},\\[10.0pt] \left(\bigotimes_{k=1}^{\nu-\left\lfloor\frac{i-1}{2}\right\rfloor}\sigma_{1}\right)\otimes\sigma_{3}\otimes\left(\bigotimes_{k=1}^{\left\lfloor\frac{i-1}{2}\right\rfloor-1}\sigma_{0}\right)&\text{if }3\leq i\leq 2\nu\text{ and }i\text{ is odd},\\[10.0pt] \sigma_{2}\otimes\left(\bigotimes_{k=1}^{\nu-1}\sigma_{0}\right)&\text{if }i=2\nu.\end{cases}$ (63)
@@
Figures 26 and 27 show the behavior of the spectral statistics for two choices of operator, $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ and $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$ for different times. In both cases, the arcsine law emerges for timescales $t\sim O(10^{1})$.
@@
Figure 26: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ for $1000$ realizations of the SYK4 Hamiltonian (60), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The number of bins in each histogram is $100$.
@@
Figure 27: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$ for $1000$ realizations of the SYK4 Hamiltonian (60), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The number of bins in each histogram is $100$.
@@
Figures 28 and 29 show the behavior of the spectral statistics for the two choices of the operator as in the SYK4 case above, $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ and $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$, for different times. In both cases, the arcsine law never emerges, at least within timescales up to order $t\sim O(10^{6})$.
@@
Figure 28: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ for $1000$ realizations of the SYK2 Hamiltonian (64), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The number of bins in each histogram is $100$. The acrsine distribution does not form, at least within timescales up to order $t\sim O(10^{6})$.
@@
Figure 29: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$ for $1000$ realizations of the SYK2 Hamiltonian (64), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The number of bins in each histogram is $100$. The arcsine distribution does not form, at least within timescales up to order $t\sim O(10^{6})$.
@@ 7 Conclusions and discussions
Strictly speaking, freeness can only be realized for operators belonging to type III von Neumann algebras, which may emerge in the thermodynamic limit, $\mathcal{D}\to\infty$, where $\mathcal{D}$ denotes the matrix rank of the operators in the algebra or, equivalently, the dimension of the Hilbert space. For finite-dimensional systems, the algebra of operators is always type I, which does not allow strict mixing. In such cases, mixed cumulants in chaotic systems do not vanish at large times but instead approach a small non-zero value that decreases with increasing $\mathcal{D}$. This leads to the notion of asymptotic freeness, where cumulants vanish only in the limit $\mathcal{D}\to\infty$.
@@
Given two deterministic matrices $A$ and $B$, and a Haar-random unitary matrix $U$, it can be shown that $A$ and $U^{\dagger}BU$ are asymptotically free in the large-$\mathcal{D}$ limit with respect to the map $\varphi_{\mathcal{D}}(\cdot)$ (19), where $\mathbb{E}$ denotes the appropriate ensemble average over the unitary $U$. Since a strongly chaotic dynamics can often be approximated by replacing the time evolution operator with a Haar-random unitary, it is natural to expect that, in realistic Hamiltonians appearing in quantum many-body physics, strong chaos would lead to $A(0)$ and $B(t)$ becoming asymptotically free with respect to the state $\varphi_{\mathcal{D}}(\cdot)$ with $\mathbb{E}$ denoting an average over possibles random parameters in the Hamiltonian; or $\varphi_{\mathcal{D}}(\cdot)={\rm Tr}\,(\cdot)/\mathcal{D}$ if the Hamiltonian does not contain any random parameters. In this case, a smoking gun of asymptotic freeness is the spectral statistics of $A(0)+B(t)$ following the free convolution prediction.
@@
As can be seen from eq. (16), the condition that defines two free operators depends on the map $\varphi(\cdot):\mathcal{A}\rightarrow\mathbb{C}$. Since, for a quantum system, this map essentially represents the state of the system, whether two quantum operators are free with respect to each other depends on this state. The primary quantity we have used in this paper as a signature of asymptotic freeness of two deterministic operators ($A$ and $B$) under time evolution, namely the spectral statistics of an operator of the form $A(0)+B(t)$, refers to a specific state of the quantum system for which the map acts as $\varphi(\cdot)=\mathbb{E}\left({\rm Tr}\,(\cdot)/\mathcal{D}\right)$. [^16] Note that the expansion of $G(z)$ (the Cauchy transformation of the eigenvalue density) in terms of the moments uses this map explicitly (see the discussion in footnote 8). In the context of random matrix ensembles, $\mathbb{E}$ refers to an average over the probability weight of the ensemble, whereas, in the present context of realistic quantum many-body systems, this is essentially an average over random parameters that might be present in the Hamiltonian. For a generic quantum system, this map would essentially represent an expectation value taken with respect to the infinite temperature thermal density matrix. Therefore, the observation that for chaotic systems, the eigenvalue density of $A(0)+B(t)$ starts to closely mimic the free convolution prediction around scrambling time is a strong indication that $A(0)$ and $B(t)$ indeed become asymptotically free under time evolution with respect to the above map.171717 We emphasize that, for a finite $\mathcal{D}$, the mixed free cumulants of $A(0)$ and $B(t)$ would not vanish at late times, but they are expected to reach a stationary residual value whose amplitude decreases as $\mathcal{D}$ is increased, indicating that $A(0)$ and $B(t)$ become free in the limit $\mathcal{D}\rightarrow\infty$.
@@
The above discussion raises the question of whether we can test asymptotic freeness with respect to some other quantum state, say a finite temperature density matrix, in a similar way. As argued above, since the free probability prediction of the eigenvalue distribution of an operator of the form $A(0)+B(t)$ makes use of a specific instance of the map $\varphi(\cdot)$, in our opinion the eigenvalue distribution of $A(0)+B(t)$ is not a suitable quantity for that purpose. However, it would be interesting to look for other measures that might be useful in that context.
@@
In the SYK4 model, the spectra of operators of the form $\psi_{i}(0)+\psi_{j}(t)$ are presented in Figures 26 and 27 for the cases $i=j=1$ and $i=1,j=5$, respectively. In both cases, the eigenvalue density converges to the arcsine law, the free probability prediction in this context, for times around $t\approx 10$. This indicates that the operators $\psi_{i}(0)$ and $\psi_{j}(t)$ become asymptotically free on this timescale. Notably, unlike the spin chain cases – where certain operators exhibit slight deviations from the free probability prediction – the SYK4 model shows no such discrepancies. This confirms the expectation that the SYK4 model is more chaotic than the spin chain models considered in this work.
The emergence of the free probability prediction in the SYK4 model at finite $N$ suggests that the model satisfies the F-property in the limit $N\to\infty$. Since the F-property implies mixing, which typically requires a type III algebra, this provides strong evidence that the corresponding algebra of observables becomes type III in this limit, in agreement with expectations from the literature Leutheusser:2021frk .
@@
In SYK2, the spectra of operators of the form $\psi_{i}(0)+\psi_{j}(t)$ are shown in Figures 28 and 29 for the cases $i=j=1$ and $i=1,j=5$, respectively. We observe that, regardless of how large the time parameter is, the arcsine law never emerges. As observed before, the same happens at the integrable point of the mixed-field Ising model.
@@
Finally, we comment on previous works that connect different notions of quantum chaos. In Cotler:2017jue , the authors establish a connection between correlators involving Haar random unitary operators and a generalized version of the spectral form factor (see Appendix D for a brief review). Specifically, they show that higher-order OTOCs involving Haar random unitaries are related to the $k$-th power of the spectral form factor, thus linking the vanishing of OTOCs with spectral statistics exhibiting random matrix behavior. In contrast, the concept of freeness between operators $A(0)$ and $B(t)$ encompasses the vanishing of all mixed free cumulants, including two-point functions, OTOCs, and higher-order out-of-time-ordered correlators. If the time evolution operator is modeled by a Haar random unitary $U$, one can show that $A(0)$ and $U^{\dagger}BU$ are asymptotically free, implying the vanishing of all mixed free cumulants. In this case, an ensemble average is taken in the definition of the map $\varphi(A)=\lim_{\mathcal{D}\rightarrow\infty}\mathbb{E}\left(\frac{{\rm Tr}\,A}{\mathcal{D}}\right)$.

### v3

@@ 6 Asymptotic Freeness in the High-Spin Mixed-Field Ising Model
$X^{(s)}_{j},Y^{(s)}_{k}]=i\frac{\sqrt{3}}{\sqrt{s(s+1)}}\delta_{jk}Z^{(s)}_{j}$, (82)
@@ 6.1 Approximate emergence of asymptotic freeness
In the spin-1 case, the fundamental operators of the theory, such as $X^{(s)}_{i}$, and their time-evolved counterparts, $X^{(s)}_{i}(t)$, have spectra characterized by a generalized Bernoulli distribution with eigenvalues $\pm\sqrt{3/2},0$. For convenience, we rescale these operators by a factor of $\sqrt{2/3}$, ensuring their spectra become $\pm 1,0$. As shown in Section 4.1 (Example 2), the free probability prediction for the spectrum of the sum of two free variables with generalized Bernoulli distribution with eigenvalues $\pm 1,0$ is given by Eq. (53). Figure 24 shows the spectrum of eigenvalues of the operator $\sqrt{2/3}(Z_{1}(0)+Z_{1}(t))$ with increasing time for 100 realizations of the Hamiltonian (80) at the chaotic point. The distribution predicted by free probability theory (53) becomes evident around $t=5$.
Refer to caption $t=0.01$ Refer to captiont=1t=1
Refer to captiont=2t=2 Refer to captiont=5t=5
@@ 6.2 Fluctuations
In this subsection, we investigate the fluctuations in the spectrum of the sum of spin-1 operators of the form $A+B(t)$. In the presence of chaotic dynamics, the operators $A$ and $B(t)$ are expected to become asymptotically free, with the distribution of their sum given by Eq. (53). The results are presented in Figures 25, 26, and 27. In all cases, we consider 100 realizations of the Hamiltonian (80) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$, and $t=50$.
@@
Figure 25: Spectral statistics of the operator $\sqrt{2/3}(Z_{1}^{(s)}+Z_{6}^{(s)}(t))$ based on 100 realizations of the Hamiltonian (80) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$ and $t=50$. The solid lines represent (a) the free probability prediction and (b) the Wigner $r$-parameter distribution for $\nu=2$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=1.999\pm 0.004$. The average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% on each edge.
@@
Figure 26: Spectral statistics of the operator $\sqrt{2/3}(Y_{1}^{(s)}+Y_{6}^{(s)}(t))$ based on 100 realizations of the Hamiltonian (80) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$ and $t=50$. The solid lines represent (a) the free probability prediction and (b) the Wigner $r$-parameter distribution for $\nu=2$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=2.000\pm 0.004$. The average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% on each edge.
@@
Figure 27: Spectral statistics of the operator $\sqrt{2/3}(X_{1}^{(s)}+X_{6}^{(s)}(t))$ based on 100 realizations of the Hamiltonian (80) at the chaotic point $(h_{x},h_{z},g)=(-1.05,0.5,0.2)$, with $s=1$, $L=6$ and $t=50$. The solid lines represent (a) the free probability prediction and (b) the Wigner $r$-parameter distribution for $\nu=2$. The $r$-parameter distribution was fitted using the form (106), yielding the Dyson index $\nu=2.000\pm 0.004$. The average $r$-parameter was found to be $\langle\tilde{r}\rangle=0.60$. The analysis focuses on the central 50% of the spectrum, excluding the outermost 25% on each edge.
@@ 7 Asymptotic Freeness in the Sachdev–Ye–Kitaev Model
[…]
@@
Figures 29 and 30 show the behavior of the spectral statistics for two choices of operator, $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ and $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$ for different times. In both cases, the arcsine law emerges for timescales $t\sim O(10^{1})$.
@@
Figure 29: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ for $1000$ realizations of the SYK4 Hamiltonian (83), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The bin size in each histogram is 0.06.
@@
Figure 30: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$ for $1000$ realizations of the SYK4 Hamiltonian (83), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The bin size in each histogram is 0.06.
@@
Figures 31 and 32 show the behavior of the spectral statistics for the two choices of the operator as in the SYK4 case above, $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ and $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$, for different times. In both cases, the arcsine law never emerges, at least within timescales up to order $t\sim O(10^{6})$.
@@
Figure 31: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{1}(t)\right)$ for $1000$ realizations of the SYK2 Hamiltonian (87), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure. The bin size in each histogram is 0.06. The acrsine distribution does not form, at least within timescales up to order $t\sim O(10^{6})$.
@@
Figure 32: Spectral statistics (eigenvalue density) of the operator $\sqrt{2}\left(\psi_{1}+\psi_{5}(t)\right)$ for $1000$ realizations of the SYK2 Hamiltonian (87), without fixing a chirality sector, with $N=16$ fermions for different times, indicated in each individual figure.The bin size in each histogram is 0.06. The arcsine distribution does not form, at least within timescales up to order $t\sim O(10^{6})$.
@@ 8 Conclusions and discussions
Strictly speaking, freeness can only be realized for operators belonging to type II or type III von Neumann algebras, which may emerge in the thermodynamic limit, $\mathcal{D}\to\infty$, where $\mathcal{D}$ denotes the matrix rank of the operators in the algebra or, equivalently, the dimension of the Hilbert space. For finite-dimensional systems, the algebra of operators is always type I, which does not allow strict mixing. In such cases, mixed free cumulants in chaotic systems do not vanish at large times but instead approach a small non-zero value that decreases with increasing $\mathcal{D}$. This leads to the notion of asymptotic freeness, where the free cumulants vanish only in the limit $\mathcal{D}\to\infty$.
@@
Given two deterministic matrices $A$ and $B$, and a Haar-random unitary matrix $U$, it can be shown that $A$ and $U^{\dagger}BU$ are asymptotically free in the large-$\mathcal{D}$ limit with respect to the map $\varphi_{\mathcal{D}}(\cdot)$ in (20), where $\mathbb{E}$ denotes the appropriate ensemble average over the unitary $U$. Since a strongly chaotic dynamics can often be approximated by replacing the time evolution operator with a Haar-random unitary, it is natural to expect that, in realistic Hamiltonians appearing in quantum many-body physics, strong chaos would lead to $A(0)$ and $B(t)$ becoming asymptotically free with respect to the state $\varphi_{\mathcal{D}}(\cdot)$, with $\mathbb{E}$ now denoting an average over possibles random parameters in the Hamiltonian; or $\varphi_{\mathcal{D}}(\cdot)={\rm Tr}\,(\cdot)/\mathcal{D}$ if the Hamiltonian does not contain any random parameters. In this case, a smoking gun of asymptotic freeness is whether the spectral statistics of $A(0)+B(t)$ follow the free convolution prediction.
@@
As can be seen from the Eq. (17), the condition that defines two free operators depends on the map $\varphi(\cdot):\mathcal{A}\rightarrow\mathbb{C}$. Since, for a quantum system, this map essentially represents the state of the system, whether two quantum operators are free with respect to each other depends on this state. The primary quantity we have used in this paper as a signature of asymptotic freeness of two deterministic operators ($A$ and $B$) under time evolution, namely the spectral statistics of an operator of the form $A(0)+B(t)$, refers to a specific state of the quantum system for which the map acts as $\varphi(\cdot)=\varphi_{\mathcal{D}}(\cdot)$. [^24] Note that the expansion of $G(z)$ (the Cauchy transformation of the eigenvalue density) in terms of the moments uses this map explicitly (see the discussion in footnote 17). In the context of random matrix ensembles, $\mathbb{E}$ refers to an average over the probability weight of the ensemble, whereas, in the present context of realistic quantum many-body systems, this is essentially an average over random parameters that might be present in the Hamiltonian. For a generic quantum system, this map would essentially represent an expectation value taken with respect to the infinite temperature thermal density matrix. Therefore, the observation that for chaotic systems, the eigenvalue density of $A(0)+B(t)$ starts to closely mimic the free convolution prediction around scrambling time is a strong indication that $A(0)$ and $B(t)$ indeed become asymptotically free under time evolution with respect to the above map.252525 We emphasize that, for a finite $\mathcal{D}$, the mixed free cumulants of $A(0)$ and $B(t)$ would not vanish at late times, but they are expected to reach a stationary residual value whose amplitude decreases as $\mathcal{D}$ is increased, indicating that $A(0)$ and $B(t)$ become free in the limit $\mathcal{D}\rightarrow\infty$.
@@
The above discussion raises the question of whether we can test asymptotic freeness with respect to some other quantum state, say a finite temperature density matrix, in a similar way. As argued above, since the free probability prediction of the eigenvalue distribution of an operator of the form $A(0)+B(t)$ makes use of a specific instance of the map $\varphi(\cdot)$, in our opinion the eigenvalue distribution of $A(0)+B(t)$ is not a suitable quantity for that purpose. However, it would be interesting to look for other measures that might be useful in that context. In appendix E, we elaborate on possible approaches to extend our results to finite temperature.
@@
In the SYK4 model, the spectra of operators of the form $\psi_{i}(0)+\psi_{j}(t)$ are presented in Figures 29 and 30 for the cases $i=j=1$ and $i=1,j=5$, respectively. In both cases, the eigenvalue density converges to the arcsine law, the free probability prediction in this context, for times around $t\approx 10$. This indicates that the operators $\psi_{i}(0)$ and $\psi_{j}(t)$ become asymptotically free on this timescale. Notably, unlike the spin chain cases – where certain operators exhibit slight deviations from the free probability prediction – the SYK4 model shows no such discrepancies. This confirms the expectation that the SYK4 model is more chaotic than the spin chain models considered in this work.
The emergence of the free probability prediction in the SYK4 model at finite $N$ suggests that the model satisfies the F-property in the limit $N\to\infty$. Since the F-property implies mixing, which typically requires a type II or III algebra, this provides strong evidence that the corresponding algebra of observables becomes type II or III in this limit, in agreement with expectations from the literature Leutheusser:2021frk .
@@
In SYK2, the spectra of operators of the form $\psi_{i}(0)+\psi_{j}(t)$ are shown in Figures 31 and 32 for the cases $i=j=1$ and $i=1,j=5$, respectively. We observe that, regardless of how large the time parameter is, the arcsine law never emerges. As observed before, the same happens at the integrable point of the mixed-field Ising model.
@@
Relation between freeness and random matrix universality
Our analysis of decorrelated ensembles in subsection 4.2, along with the emergence of freeness near the integrable point in spin chains, suggests that freeness can arise even without random matrix universality. In such cases, however, it appears in an operator-dependent manner – some operators become free from their past versions or other past operators, while others, particularly those linked to conserved charges, do not. Since freeness essentially captures the notion of strong operator growth or scrambling, this suggests that while random matrix universality implies scrambling, the converse is not true: scrambling can occur without random matrix behavior, albeit in an operator-dependent way. This is consistent with the findings of Dowling:2023hqc , which state that scrambling is a necessary but not sufficient condition for random matrix chaos. Here we emphasize that by scrambling we mean the vanishing of OTOCs and their higher-order counterparts, not any possible exponential behavior.
@@
Finally, we comment on previous works that connect different notions of quantum chaos. In Cotler:2017jue , the authors establish a connection between correlators involving Haar random unitary operators and a generalized version of the spectral form factor (see Appendix D for a brief review). Specifically, they show that higher-order OTOCs involving Haar random unitaries are related to the $k$-th power of the spectral form factor, thus linking the vanishing of OTOCs with spectral statistics exhibiting random matrix behavior. In contrast, the concept of freeness between operators $A(0)$ and $B(t)$ encompasses the vanishing of all mixed free cumulants, implying the vanishing of two-point functions, OTOCs, and higher-order out-of-time-ordered correlators for traceless operators. If the time evolution operator is modeled by a Haar random unitary $U$, one can show that $A(0)$ and $U^{\dagger}BU$ are asymptotically free, implying the vanishing of all mixed free cumulants. In this case, an ensemble average is taken in the definition of the map $\varphi(A)=\lim_{\mathcal{D}\rightarrow\infty}\mathbb{E}\left(\frac{{\rm Tr}\,A}{\mathcal{D}}\right)$.